OpenStudy (anonymous):

If g(x) = 3x - 5, then g -1(x)=

OpenStudy (anonymous):

Switch the X and Y, then solve for y xD So x=3y-5.

OpenStudy (anonymous):

x+5=3y......... Divide by 3.... x+5/3=y

OpenStudy (anonymous):

\[\frac{ x+5 }{ 3}=y\]

OpenStudy (anonymous):

g-1(x) = inverse g(x) = 3x-5 Let g(x) = y y = 3x-5 x = 3y-5 x+5 = 3y (x+5)/3 = y therefore, g(x)^-1 = (x+5)/3